OpenStudy (anonymous):

Which function below is the inverse of f(x) = x2 − 25?

OpenStudy (anonymous):

it is \[x^2 ?\]

geerky42 (geerky42):

very likely

OpenStudy (anonymous):

x squared over 25 ±square root of the quantity x plus 25 ±5 times the square root of x 1 over quantity x squared minus 25

OpenStudy (anonymous):

@Yahoo!

OpenStudy (anonymous):

put y =f(x) y = x^2 - 25 x^2 = y+25 x = sqrt(y+25) or x = -sqrt(y+25) so f ^-1 (x) = sqrt(x+25) or -sqrt(x+25)

OpenStudy (anonymous):

im lost @Yahoo!

OpenStudy (anonymous):

LOL....tell me where u lost

OpenStudy (anonymous):

x =sqrt(x+25)

OpenStudy (anonymous):

if x^2 = 9 then wat is x?

OpenStudy (anonymous):

ooooh i got it now lol thanks

OpenStudy (anonymous):

Welcome..:)